### v1

# Do you monitor CI Practices? I don’t know. You tell me: A case study

Abstract.
Background: In this paper we seek for understand the benefits and challenges of monitoring CI practices in day-to-day software development. Aims: We aim to evaluate the impact of monitoring seven CI practices in a real-world scenario on three public organizations in Brazil. Method: We first developed a CI practices monitoring suite tool and conducted a multiple-case study applying a mixed-methods strategy, combining surveys, interviews, log data, and mining data from CI services. Results: We verified organization’ interest in monitoring CI practices. Monitoring provided an overview of the organization’s CI status, not covered by other tools, motivated constant improvement in these practices, a perception of software quality, improve the communication and and it is easy to adopt. Conclusions: We recommend that companies adopt monitoring of CI practices and that CI services integrate this monitoring into their dashboards.
CI Practices, Monitoring, Case Study, CI Maturity
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Software performanceccs: Software and its engineering Software reliability
1. Introduction
Continuous Integration (CI) is a software development practice where team members integrate their work frequently (Fowler, 2006). The integration process involves an automated build step, which includes compiling the code and running automated tests. This ensures the detection of integration errors as quickly as possible. Despite of extensive research into the benefits and costs of Continuous Integration  (Laukkanen et al., 2015; Widder et al., 2019; Hilton et al., 2017; Cassee et al., 2020; Hilton et al., 2016; Ståhl and Bosch, 2013a; Elazhary et al., 2021; Vasilescu et al., 2015; Bernardo et al., 2018; Ghaleb et al., 2019; Bach et al., 2017). Most of studies focused in miner repositories and study the overall impact of CI. There is a lack of studies specifically addressing the monitoring CI practices. Do developers monitor CI practices properly, and would they like to monitor? What benefits do they receive from this monitoring? These are some of the questions we aim to answer in this paper.
In their Systematic Literature Review, Eliezio et al. (Soares et al., 2021) identified a lack of criteria for defining the use of CI. They found that 42.5% of the primary studies did not apply or establish any criteria to determine whether a project uses CI. Among those that did use criteria, 56.25% of the studies relied on only one criterion. The most common criterion applied was the adoption of a CI service, such as Travis CI. In the same direction, other studies (Felidré et al., 2019) have already demonstrated that some CI projects do not follow all CI practices. That study found that: (i) 60% of the projects had infrequent commits; (ii) 85% of the projects had at least one broken build that took a long time to be fixed; and (iii) most projects had builds that executed for more than 10 minutes. A fact known as CI THEATER.
In this context, this paper aims to understand the benefits of using and monitoring CI practices by assessing their utility in real-world software development projects. To achieve this goal, we conducted a multiple case study with mixed methods, combining survey, interviews and access log analysis. For each project, we applied an initial survey to understand the level of CI maturity and the team’s satisfaction with CI. This allowed us to better compare and understand the impact of monitoring CI practices on projects at the end of the study. Additionally, we developed a tool with an extensive suite of metrics that automatically calculates the values of seven CI practices, enabling developers to continually have access to these metrics and receive weekly alerts about CI practice evolution. The CI practices are describe bellow:
  • Commit Per Weekday/Commit Activity  (Felidré et al., 2019): Mean of the absolute number of commits according to the week day of the analyzed period.
  • Coverage  (Felidré et al., 2019): It measures which parts of a program are executed when running the tests. Represents the percentage of the program covered by tests.
  • Build Duration  (Felidré et al., 2019): It measures the duration of the build (build finished at timestamp - build started at timestamp).
  • Build Activity/Build-frequency  (Ståhl and Bosch, 2014): It is a unit interval (i.e., a closed interval [0,1]) representing the rate of builds across days, i.e, if builds were made every day in the period of CI, the value would be 1. If builds were made in half of the days, the value would be 0.5. If there were no builds, the value would be 0.
  • Build Health/Build Quality  (Soares et al., 2021): It is a unit interval representing the rate of build failures across days. If there were build failures every day, the value would be 0. if there were no build failures, the value would be 1.
  • Time to Fix a Broken Build  (Felidré et al., 2019): It consists of the median time in a period that builds remained broken. When a build breaks, we compute the time in seconds until the build returns to the “passed” status. If the CI period ends and the build did not return to the “passed” status, we consider the time since it was broken until the end of the period. When a period has no broken builds, the value would be 0.
  • Comments Per Change  (Thompson, 2017; Ståhl and Bosch, 2013b): Mean of the number of comments grouped by Merge Request or Issue. This metric seeks to measure the level of communication between the team.
Over a period of 8 weeks, we conducted a series of interviews with team members to gain insights into the benefits and challenges of monitoring these CI practices. We applied this suite of metrics in three software development projects from different public Brazilian organizations. By collecting metrics from CI services, access log, surveys and interviewing organizational members, we answered the following research questions:
  • RQ1: Are the developers satisfied with the project’s CI maturity levels? By surveying and interviewing members of the project, we assess whether developers are satisfied with the current level of CI in their project. This exploratory research question gives us insight into the need to monitor CI practices. Our findings indicate that $23\frac${2}{3}divide start_ARG 2 end_ARG start_ARG 3 end_ARG of the projects were satisfied with the CI maturity level before the case study, although they did not utilize any metric-driven assessment to evaluate CI maturity. Nevertheless, despite initial satisfaction, CI monitoring helped participants to notice several aspects of CI that needed to be improved.
  • RQ2: What benefits does CI monitoring bring? By interviewing members of the project, we collected team members’ perceptions regarding the benefits of monitoring CI practices. Our findings show that participants identified several benefits of CI monitoring, including its ability to provide an overview of the CI stage, highlight problems related to CI, and serve as a reminder and motivator for CI improvement, among other advantages.
  • RQ3: What are the challenges/disadvantages of monitoring CI practices on a daily basis? By interviewing members of the project, we collected team members’ perceptions regarding challenges faced in monitoring CI practices. There were no significant challenges directly related to monitoring CI practices and the use of the monitoring tool. External problems, such as deadline management, instability in the CI environment, and a lack of emphasis on the importance of CI, were the main challenges faced by projects.
  • RQ4: Are developers interested in using the tool? By collecting the log of access to dashboard of our suite of metrics, without the knowledge of the participants, we assessed how frequently the tool was used during the case study by the development team, thus measuring real interest in monitoring CI practices. The record of access to the dashboard screen allows us to conclude that there was considerable access to the tool during the case study, even outside the interview days. Mainly by companies that already had a certain level of CI maturity.
  • RQ5: What was the evolution of CI practices during the use of the tool? By comparing the evolution of CI practices throughout the study period, we gain insights into the role of monitoring in this evolution. We noticed a notable enhancement in CI practices with one project, but for the other two projects, due to the instability of the CI environment and the break for the end-of-year holiday, we cannot perceive a constant improvement.
2. Continuous Integration Practices Monitoring
Wang et. al. (Wang et al., 2022) presented an empirical study to observe the impact of test automation maturity on product quality, test automation effort, and release cycles in the Continuous Integration (CI) context of open-source projects. They carried out a test automation maturity survey with the main contributors of 149 open-source Java projects. “Test automation maturity” was quantified as the total score from all the maturity-related questions in their survey. Their findings indicated that higher levels of test automation maturity were positively correlated with improved product quality (p-value=0.000624) and shorter release cycles (p-value=0.01891).
Sallin et. al.  (Sallin et al., 2021) conducted a comprehensive literature review to assess the feasibility of automatically measuring the four DORA Metrics (GmbH, 2023). They discovered that no scientific literature had specifically investigated the automatic measurement of these metrics. Nevertheless, their examination of grey literature revealed 16 papers discussing various aspects of automatic measurement. Additionally, they conducted a case study involving a Scrum team comprising ten individuals to evaluate the benefits of automating the measurement of DORA metrics.
Elazhary et. al. (Elazhary et al., 2021) investigated the extent to which three software organizations implement ten Continuous Integration (CI) practices defined by (Fowler, 2006). They explored the benefits these practices bring and the challenges encountered during their implementation. This inquiry was conducted through a multiple-case study with mixed methods, focusing on three small to medium-sized software-as-a-service organizations. The study revealed that four practices exhibited significant differences in adoption between organizations: (i) For “Maintain a single source repository” practice: dealing with merge conflicts had a higher priority than the unclear benefit of using a single source repository, (ii) For “Make the build self-testing” practice: the perceived work required for integration tests outweighs the perceived benefits, (iii) For “Keep the build fast” practice: doing comprehensive testing outweighs having shorter builds), (iv) For “Automate deployment” practice: enhancing security by enforcing deployment privileges for some organizations outweighs a sense of change ownership.
This paper aligns with the previously cited works. While Wang et. al. (Wang et al., 2022) concentrated on a single CI practice, specifically test automation, this paper takes a broader, more exploratory approach by evaluating a larger group of practices. In contrast with the Sallin et. al.  (Sallin et al., 2021) work, this paper focus on scope of CI practices. Furthermore, the paper aimed to evaluate the advantages of monitoring these metrics, extending beyond the focus solely on automation. Perhaps the work most aligned with ours is the work proposed by Elazhary et. al.(Elazhary et al., 2021). However, our analysis focuses on different CI practices. Additionally, a dashboard tool was developed as part of this paper, which automates the collection and visualization of the CI practices studied, thereby facilitating the adoption of monitoring CI practices. In our study, we implemented a metric suite tool that automatically collects and delivers 7 CI practice values to developers. This represents a significant advantage, particularly for the organization undergoing the case study. Our primary focus lies in assessing the benefits and challenges associated with the monitoring of CI practices over time. In contrast, Elazhary et. al. (Elazhary et al., 2021) emphasis is on exploring the benefits and challenges of implementing CI practices.
3. Methodology
We conducted a multiple-case study (Stake, 2013) on monitoring CI practices with three software projects from different organizations. The main inclusion criteria for organizations in this study were: (i) location – we opportunistically chose organizations where we could interact with the developers and personally visit the organization to give more confidence in analyzing their data; (ii) availability – organizations had to provide a basic infrastructure to run our tool and allocate some of their developers time for interviews and validation of our findings; and (iii) the organization should have at least one project that should be using a CI service and generating regular builds.
The characteristics of selected organization are:
  • Department of IT at a Public University (Organization A): Organization A is a public educational institution, specifically a department of Information Technology (IT) within a prominent public university in Brazil. It encompasses a wide range of academic and administrative IT projects, such as: a job vacancy system, “technology park” management system, access control system, website, etc. The university is known for its commitment to research and technological advancements, making it an ideal candidate for studying CI practices.
  • Software Project at a University with a Global Technology Company (Organization B): Organization B represents a unique collaboration Organization Between a Brazilian university and a global technology company. This software project focuses on cutting-edge technology and innovation in 5G networks. The involvement with a global technology company brings a distinct perspective on CI practices and the integration of industry and academia.
  • State House of Representatives (Organization C): Organization C represents a branch of the state government responsible for legislative activities. As a public institution, the State House of Representatives plays a vital role in shaping legislation and policies at the state level in Brazil. The development team work on developing the project’s management software, such as: human resources department system, website, among others. The inclusion of this organization allows for a diverse perspective on CI practices, different from the academic environment.
Because of the ease of obtaining access authorizations, Organization A was the first to initiate the interviews of case study, starting the interviews in November 2023 and concluding them in December 2023. We took advantage of this to use this organization as a pilot for our case study, as suggested by best practices (Yin, 2009). Organization B and C started the case study interviews in mid-December 2023 and finished it in early February 2024. We collected CI practice data from October 1, 2023, until the end of the interview period. This resulted in the evolution of practices being tracked over a period of 3 months for Organization A and 4 and a half months for Organizations B and C.
Coincidentally, all three organizations participating in the case study utilized the same CI service, Gitlab. They employed Gitlab On-Premise, installed on the organizations’ Intranet, requiring specific access requests for each entity. To gather CI practices, we mined the build history of projects in the Gitlab CI service. Additionally, we collected the coverage metric from the SonarQube tool whenever the project used it (Organizations A and B utilized this tool, while Organization C did not generate these metric). We selected key members of the development team as study participants, considering developers with experience and basic knowledge of CI and the company’s development environment. Figure 1 provides an overview of all the steps involved in our data collection plan.
Figure 1. Data collection plan
The data collection plan for this case study involves eight-step approach. We start with an initial presentation to introduce the case study’s objectives, importance, needs and potential benefits for projects that participate in it. This step was important to convince organizations to participate in the case study. Following the initial presentation, the project’s representative are asked to sign privacy declarations to ensure data confidentiality and obtaining permission from the projects to proceed with the study. We apply an initial survey (explained in detail in Section 3.1) to measure the project current CI maturity level perception for all study participants. Subsequently, training sessions are conducted to familiarize the teams with the developed metric suite tool for CI monitoring. The tool is then installed in the Intranet of organization, on a server provided by them. This makes it easier for participants to access our tool, which was available 24 hours a day in their work environment. The suite metric tool is then configured to start the weekly collection of CI practices. Finally, we perform weekly interviews with the development teams (explained in detail in Section 3.2) to capture real-time insights to measure the impact of CI practices monitoring.
In conducting the case study, we adhered to a systematic and structured methodology based on (Tellis et al., 1997). 1) The first phase involved the design of a comprehensive case study protocol. This protocol, should determinate the required skills for case study and should be developed and reviewed. 2) The second phase centered on the actual execution of the case study. Adequate preparation for data collection, installation of CI practices monitoring suite and the distribution of questionnaires to relevant participants within the selected projects to assess the perception of current CI maturity in the project. Subsequently, interviews were conducted weekly to gather qualitative insights focused on answering the study’s research questions. 3) The third phase focused on the meticulous analysis of the gathered evidence. 4) The fourth phase involved the synthesis of conclusions, recommendations, and implications based on the evidence obtained through the case study. We also following the best practices defined by (Yin, 2009): 1) Conducting a pilot case study (Organization A), 2) training for a specific case study, 3) screening candidate cases, 4) Ask good questions, 5) Be a good “listener”, 6) Stay adaptive, 7) Have a firm grasp of the issues being studied and 8) Avoid biases by being sensitive to contrary evidence.
To encourage participation in the study, we also outline the benefits for projects participating in the study, they are: (i) the project will be able to have a overview vision of how its CI maturity process, (ii) The monitoring suite tool, along with its source code, was made available permanently and free of charge to the project, and (iii), at the end of the study, 10 more metrics will be unlocked, including Continuous Delivery and DevOps metrics (DORA metrics  [^1] https://cloud.google.com/blog/products/devops-sre/using-the-four-keys-to-measure-your-devops-performance) .
3.1. Survey
To gain an understanding of the project’s CI maturity before implementing the case study, we requested each developer to answer a preliminary survey. This initial survey aided in comprehending the CI state within each project prior to the case study, assessing satisfaction with CI adoption, and identifying any metrics used to measure CI. This preliminary insight facilitated a more comprehensive evaluation of the case study’s impact on the evolution of CI in each analyzed project. Furthermore, the survey responses helped us answer RQ1.
The survey consisted of 15 questions, 9 of which were open-ended. To give respondents as much flexibility as possible, 9 questions were mandatory (6 of which were close-ended while 3 was open-ended), and 6 questions were optional (all open-ended). The survey was divided into four sections: 1) Introduction, 2) Personal information, 3) Team Organization And Collaboration, and 4) Diagnosis of the project’s CI maturity. A full copy of the survey is available at https://zenodo.org/records/11106110.
In the “Introduction” section of the survey, we provide context for the study, introduce Continuous Integration concept, and outline the survey’s objectives. We also specify the estimated time required for completion and assure participants that their responses will be used solely for research purposes and published anonymously. The “Personal Information” section assesses participants’ knowledge of software development and Continuous Integration, ensuring they have the necessary qualifications for the case study. The “Team Organization and Collaboration” section explores the organization’s CI environment, development processes, and tools used. Finally, the “Diagnosis of the Project’s CI Maturity” section investigates how the organization evaluates the CI maturity of its projects, the metrics employed, and overall satisfaction with the CI adoption.
3.2. Interviews
We conduct semi-structured interviews (Dearnley, 2005) with the participants to obtain insights about monitoring CI practices of analyzed projects. We defined a set of initial questions, although they could vary depending on the participants’ responses during the interviews. The interview questions were used primarily to address the research questions RQ2 and RQ3 and are detailed in the case study protocol. We follow the guidance provided by (Hancock et al., 2021). We also follow the best practices before the Interview: 1) Identifying Key Participants, 2) Developing an Interview Guide/Protocol, 3) Ensuring a Private and Neutral Environment, 4) Recording the Interview Data, 5) Adhering to Legal and Ethical Requirements. Conducting the Interview we: 1) Obtaining Consent and Ensuring Anonymity, 2) Reviewing the Interview’s Purpose and Duration, 3) Listening Actively. We labeled each participant of case study to protect anonymity, Table 1 shows study participants.
Table 1. Case Study Participants
Label Role Project
P1 Software Developer A
P2
Software Developer;
Software Testing Analyst;
Application Architect
B
P3 Researcher B
P4
Software developer;
5G networks Researcher
B
P5 Software Developer C
3.3. Thematic Analysis of Interview Data
Following a similar approach to (Elazhary et al., 2021), we transcribed the interviews of all five participants. We assigned a label to participants’ statements in the interviews that sought to express, in a simple and summarized way, the participant’s feelings regarding monitoring CI practices. We grouped common themes and created a network graph to represent them. From there, we carried out an analysis to extract the main findings of these themes. We sent each participant’s final themes so they could review them and express whether they agreed with the extracted themes.
3.4. Study Replications
We have made the documentation generated by this case study available to facilitate future replications at:
https://zenodo.org/records/11106110. We provide the use case protocol (in English), the tool source code, as well as an example of the survey about the project’s CI maturity. Additionally, the values of CI practices throughout the case study and interview transcriptions are included anonymously (the latter three in Portuguese).
4. Results
4.1. Thematic Analysis
Figure 2 shows the main themes that emerge from the interviews about monitoring CI practices with the participants. Each theme summarizes an idea expressed by the participants during the interviews. Each theme has a number that represents the number of mentions expressed by participants. The green nodes are themes that have a positive aspect of monitoring CI practices, while the red nodes represent negative aspects pointed out by interview participants.
Figure 2. Interviews Themes
The primary positive theme that emerged when monitoring CI practices was the “Desire for Improvements” (28 mentions) which illustrates the participant’s desire to improve some of the CI practices within the project (i.e., the participant expresses the desire to improve certain practices in the future). For instance, participant P1 expressed this sentiment by stating: “We had already been doing this testing issue before, but it certainly reinforced the need to do it as quickly as possible.”. The second most cited theme was “Highlights the Problem” (23 mentions) indicating the discovery of previously unnoticed issues by the development team through the aid of the monitoring tool, participant P2 articulated this experience, saying: “In fact, I found it strange that the coverage dropped, I thought it had already been resolved, there is something there that is being executed and resetting the data.”. “Ease of Monitoring” was the third most cited theme (18 mentions), this theme reflects the ease that developers found during the case study to monitor CI practices using our tool, participant P4 commented: “No, no, it’s not expensive. It was very intuitive”. Monitoring CI practices “Reminders and Motivators” participant of case study (16 mentions) as the tool constantly highlighted issues to developers, motivating them to improve the CI process, participant P3 remarked: “… but I think that seeing a very prominent percentage gets attention and helps convince other developers to write tests”.
Monitoring CI practices also provides an “Overview of CI State” for the project (12 mentions) enabling developers to view all CI-related aspects in one place, as participant P2 described: “… Precisely the tool summarizes several things that perhaps we need to use several tools to get to this information.”. Additionally, monitoring fosters “Encouraging Communication” among team members (12 mentions), as mention by participant P5: “I think it’s interesting to know what the practices are like. Just the fact that we talked somewhat superficially about commits would already be a benefit.”. Moreover, it promotes “CI Improvements”(9 mentions) as participant P1 emphasized: “I am more focused on developing the features that are missing and working on test coverage, which has the most impact for us. I just uploaded some testing issues that will improve the coverage”. There were also other themes mentioned to a lesser extent, such as: “Impact on Project Quality” (4 mentions), “Encouraging Colaboration” (1 mentions),“Gave a Direction” (1 mentions), and “Encouraging Changes” (1 mentions).
The primary negative theme in monitoring CI practices is “Challenges in Monitoring”. It emerged from difficulties using the tool or suggestions for improvement indicated by participants during the case study. Most of these issues were related to integrating the tool with other CI tools already used by the team, aiming to facilitate the CI practices monitoring. For example, participant P3 said: “If there was something like a link in the Gitlab interface as Sonarqube has. I could click and be redirected to the tool … ”, participant P5 comment: “I don’t know if the tool already has some type of integration with communication tools, could be a integration with the internal chat that we use for communication … ”. Integration with other tools is beyond the initial scope of this study, but it expresses the need for CI tools like GitLab to integrate monitoring of CI practices into their CI pipelines.
The “Time Management Challenges” theme is mainly related to the difficulty of the team implementing improvements in CI practices due to a lack of time dedicated to this task. This is a common complaint, as developers usually have other tasks and cannot be dedicated exclusively to improving CI practices for their projects. Overall, the monitoring tool proved to be easy and intuitive. Monitoring CI practices has not proven costly for participants. Participant P4 comment: “I think the tool is very intuitive, very didactic …”. Participant P1 said: “It has no major impact on configuration, doesn’t generate extra work, and adds significant value to this aspect of CI practices.”. Participant P5 appreciated the alerts sent by the tool: “I noticed that you can receive it by email. I found the feedback from the emails interesting.”
4.2. RQ1: Were the developers satisfied with the project’s CI maturity levels?
Organization A currently characterizes itself as operating at an INTERMEDIATE level of CI adoption. The classification at this level did not involve utilizing any specific CI practice. However, there is a recognized need for a strategic shift towards incorporating metrics into their CI process. The team believes that to start using metrics, it is necessary to carry out planning to define which CI practices would be appropriate for their context and how this would be implemented and used. When questioned about their satisfaction with the current CI maturity level, they said they were satisfied with the company’s CI level before using the monitoring suite tool. Furthermore, there is a perception that adopting automated database migration techniques, eliminating manual database upgrade routines, and incorporating fully automatic acceptance testing would significantly contribute to enhancing the project’s CI maturity. However, during the interviews, after starting to monitor CI practices, participant P1 expressed the need to focus more on automated testing: “In general, since the beginning of the project, I think so, but I am still not completely satisfied, mainly the coverage has not yet reached the reference value.”. Clearly, CI monitoring, with the constant sending of alerts, helped participants realize that there were additional areas for improvement beyond those mentioned in the initial survey. As participant P1 reported: “At the beginning, I didn’t even have a perspective on how it would be. The tool, like the study, also emphasizes the importance of providing an accurate idea of how the CI is performing, enabling individuals to understand the metrics and reference values”.
Organization B identifies itself as operating at an INTERMEDIATE level of CI maturity. The classification at this level is not determined by specific CI practices but rather by the successful automation of critical processes: the build process, automated testing, automatic generation of Docker images, and the deployment of new versions in a dedicated test/approval environment. Despite the absence of metric-driven assessment, participants express contentment with the current state of CI within the company. Nevertheless, they envision further improvements to enhance CI effectiveness. Participants believe that greater code coverage pertaining to the performed tests, incorporating tools focused on code security, and utilizing Flyway to automate database version evolution would contribute significantly to elevating the overall maturity and efficacy of the project’s CI practices. However, CI monitoring helped them realize that there were still several aspects to be improved, as participant P2 described: “It’s something we’re still maturing, we started from scratch. I had little knowledge but never learned how to install, deploy, or build a pipeline. Then, one thing ended up helping the other, seeing these metrics and what needed improvement in our projects”.
The Organization C characterizes its current level of CI maturity as a BASE level. This self-assessment lacks the incorporation of specific CI practices to precisely delineate the extent of adoption. Despite this, the case study participant expresses a notable dissatisfaction with the existing level of CI within the company, emphasizing that CI is still in its initial stages of implementation.
Our results reveal that, excluding Organization C, which is still in the initial stages of CI adoption and conducts most processes manually, the other projects were satisfied with the CI maturity level before the case study. Although they did not utilize any metric-driven assessment to evaluate CI maturity, they considered automating builds and writing tests as the criteria for defining the CI maturity of projects. However, despite initial satisfaction, CI monitoring helped participants to notice several aspects of CI that needed to be improved in their projects.
4.3. RQ2: What benefits does CI monitoring bring?
In Organization A, one of the most important benefits demonstrated by CI monitoring was its capacity to give an overview of the CI stage of a project. This feature is not usually directly implemented in CI services, as Participant P1 commented: “It brings precisely this visibility of those aspects that can be considered a problem. As I was talking about this coverage, we already have this information on the Sonarqube. But the other ones are new aspects that bring a more qualitative view of the CI practices we did not have in other tools … In GitLab, you have a view of the pipeline but do not have an aggregation like in this dashboard, in statistical terms.”. Monitoring CI practices also proved to be easy to use and without a high cost for Organization A. Participant P1 explained: “ … because it is easy to use, it brings value, even if the person does not use it very often. It has notifications so that the person is always there to know what the values are like. So, in short, it has no major impact on configuration, doesn’t generate extra work, and adds significant value to this aspect of CI practices”
In Organization B, monitoring CI practices greatly sparked communication and the desire for improvements. Regarding communication, Participant P4 highlighted: “yes undoubtedly, we started to hold internal conversations about how to interact more efficiently, we also added revisions for each commit we applied, this was not a common practice…”. Participant P3 commented on the desire to improve: “The red images (in dashboard) of the build duration caught my attention. Maybe this sparked my interest in investigating and reducing the dependencies used” It also made developers notice overlooked problems. Participant P2 describes: “It helped to highlight… The coverage is evident; it was interesting because it showed that there was coverage data, and suddenly, it disappeared. I think it’s helping.”. Participant P3 also said: “I think it … pointing out that we are having problems with the code coverage metric. Regarding the build, we already felt that the build took a long time and that the VMs were slow, but not that much, more than 8 hours of build time. We have tests, tests that involve the instantiation of other modules in VMs, …., but almost 10h is a very high value.”
In Organization C, Participant P5 stated that monitoring CI practices improve them by emphasizing: “Specifically, this concern of keeping builds always working, for example. The number of commits, worrying about always committing.”. It highlights the problems: “Having this awareness that the number of commits is important in the project. If there is no significant quantity, we may make commits with a lot of information at once. It helps to cover different aspects”. Monitoring CI practices also provides an overview of CI stage of a project, as commented by Participant P5: “It brings more confidence related to some CI needs, for sure. It provided a better view of this process.”. It also encourages communication: “We even had a conversation to discuss these metrics values.”.
Some benefits were observed for the different projects, such as: 1) The reminders and motivation provided to developers were among the most beneficial points highlighted by participants. Participant P1 said: “… it is another thing to be constantly reminded that you have to do it here.”. Monitoring CI practices is not a priority feature in other tools as Participant P2 claims: “I found that the tool consolidates a lot of information in one place that sometimes we have to look for on several Sonarqube screens. It helped a lot. It synthesized it in one place, which I think is cool.”. Finally, the functionality of constantly sending alerts to participants was remembered as a very important aspect of monitoring. Participant P1 reported: “Now with the email alert that is sent weekly. It keeps this concern with metrics at the forefront of my mind. Otherwise, if you’re focused on other project problems, it’s easy to leave it aside”.
Our findings show that the most important benefits demonstrated by CI monitoring were its capacity to give an overview of the CI stage, highlight CI-related problems, awaken a desire for improvements, improve the quality of CI adoption, and increase communication within the team. CI monitoring also proved to be not costly to apply.
4.4. RQ3: What are the challenges/disadvantages for using it on a daily basis?
One of the biggest challenges reported in Organization A to use the monitoring of CI practices daily was not directly related to monitoring CI practices but to improving the CI maturity from the results displayed by the monitoring tool, mainly due to time issues. Monitoring CI practices helps highlight the project’s lack of testing, but there were obstacles to improving these practices, as revealed by Participant P1: “To generate a new project, we end up having to do a presentation during the sprint and sometimes I get a little worried about not having time to present what has been completed if I’m going to do all the tests, then we end up having to fit the tests into the free time we have in the sprint”
When asked about the difficulties, Organization B also pointed out that the time dedicated to other development tasks was the main difficulty in monitoring and improving CI practices. Participant P2 said: “… We have very long builds with this build duration in another module, this one has a way to get around it, I just have not had the time yet.” and Participant P4: “… because the priority is the documentation that needs to be delivered by the end of the year”. The lack of integration with other CI tools was a problem also indicated as a difficulty in monitoring Organization B. Participant P3 said: “If there was something like a link in the Gitlab interface as Sonarqube has. I could click and be redirected to the tool … ” and Participant P4 mention: “maybe an idea, if you can find some way to add a plugin to Gitlab … ”
In Organization C, relevant challenges in monitoring CI practices were also not reported. Legacy projects and the internal culture of the project were the main reasons that made improving CI practices, as shown by the tool, challenging. Participant P5 described the difficulty of adopting CI due to the characteristics of the projects, for example, in legacy systems: “It’s because it uses a legacy system, but it’s just a different way of working. The person who developed a large part of the project was a DBA. He worked directly with the tables, and in the end, it was complicated.” Participant P5 also suggested that better interaction with other tools could facilitate CI monitoring: “I don’t know if the tool already has some type of integration with communication tools, could be a integration with the internal chat that we use for communication … ”
During the case study, we also observed additional issues that hinder the most appropriate CI adoption. In Organization A, the pressure for fast deliveries leads to underestimated deadlines, causing certain CI practices, such as writing automated tests, to be frequently postponed or ignored. Delayed testing may detect integration issues only when different parts of the system are combined, leading to significant delays and increased complexity in remediation. Organization B faces challenges due to a complex development environment, reflected in the instability of the CI environment, hindering any substantial improvement in the project’s CI maturity. Organization C needs to shift its project mindset to realize the full benefits of CI adoption. It was noted that other demands take priority, and developers are not actively encouraged to enhance CI maturity.
Overall, there were no big challenges directly related to monitoring CI practices. The main challenges reported were related to improving CI practices due to a lack of time for developers who are required by other project tasks or due to the internal culture of the project. Some complaints were made about the tool’s usability and the lack of integration of the monitoring suite tool (graphical interface) with CI services. Pressure for fast deliveries, complexity and instability in the CI environment, and a lack of emphasis on the importance of CI were identified as problems during monitoring, hindering the improvement of CI maturity.
4.5. RQ4: Are developers interested in using the tool?
Figure 3 shows the access data of the three organizations analyzed in the case study. The red bars represent tool access on the days when interviews were conducted. The blue bars represent access on days without interviews. The red crossed-out bars represent days when interviews took place, but no access was recorded in the tool.
Figure 3. (A) Number of Access to the monitoring tool of Organization A during the case study. (B) Number of Access to the monitoring tool of Organization B during the case study. (C) Number of Access to the monitoring tool of Organization C during the case study.
In Organization A, during the case study, we observed a total of 43 accesses to the tool dashboard of 2 different developers. Notably, the primary access points coincided with interview dates, as indicated by red bars on the timeline. These interview-centric peaks suggest heightened activity during those periods, potentially linked to collaborative discussions or focused on project evaluations. However, it is noteworthy that many accesses occurred outside the interview dates, indicating ongoing engagement and utilization of the tool beyond the case study evaluation periods. This broader pattern of access underscores the continuous and diverse interactions with the monitoring tool, emphasizing its relevance throughout the case study period.
In Organization B, we observed 31 accesses to the tool dashboard by two different participants. Although 3 participants participated in the case study, one did not access the tool. We notice many accesses outside the interview dates (blue bars), indicating interest in monitoring CI practices. There is a sharp decline in access to the monitoring tool after December 22, 2023. This decline can be attributed to two factors: (i) the project going on a holiday break; and (ii) one of the participants taking vacation in the last two weeks of the case study. Another consideration is that the project was in a phase with few changes. Combined with receiving the values of CI practices via email, this caused a decrease in interest in accessing the tool interface, as explained by participant P3: “I received the email and the issue of code coverage appeared, as I already knew about code coverage, I did not pay much attention to the tool because I thought that perhaps it would be more of a priority to look at the issue of the pipeline, … , for the tool to return the collection of the coverage metric correctly.”
In the last project, Organization C, there was the more reduced access to the monitoring tool, with only 7 accesses to the tool dashboard by 1 participant. This could be justified by the company’s cultural resistance to CI adoption. Additionally, the project had a holiday break at the end of the year, which is reflected in the access curve dropping and rising again when Organization Activities resumed. Another observation is that only one participant had access to the monitoring tool during the case study. This could potentially hinder the shift in the company’s culture regarding CI.
The registration of accesses to the dashboard screen allows us to conclude that projects A and B, which had defined themselves as having an INTERMEDIATE CI maturity level, showed greater interest in accessing the tool. Organization C, which has a BASIC CI maturity, exhibited a reduced interest in using the tool.
4.6. RQ5: What was the evolution of CI practices during the use of the tool?
Figure 4 highlights the evolution of CI practices for the three organizations during the case study.
Figure 4. (A) Organization A CI Practices Evolution during the case study. (B) Organization B CI Practices Evolution during the case study. (C) Organization C CI Practices Evolution during the case study.
In Organization A, since the initiation of our monitoring efforts, all CI practices have demonstrated considerable improvement. Notably, the “Time to Fix a Broken Build” experienced a significant reduction, indicative of enhanced efficiency in addressing issues promptly. As a consequence, the “Build Health” has increased significantly. The improvement in test coverage is equally relevant, and interviews revealed that monitoring practices played a crucial role in driving this positive trend. Participant P1 said: “I think the evaluation is good for us to have more detailed information, so we can make changes to the project, where necessary. We already knew about the coverage problem from the beginning, but it became very evident. With this information, we began to act to resolve this problem.”. The “Commits per Day” practice dropped in the middle of the case study, followed by a strong recovery. We associate it to the departure of a team member and, subsequently, the entry of new members into the development team. The same event of change in the development team may have influenced the “Comments per Change” practice, which increased during the case study.
Organization B presented a very unstable CI environment due to the complex integration environment with other applications, reflected in the CI practices. The code coverage mostly remained zero despite the project having automated tests. The build time fluctuated greatly between normal weeks and weeks with very high peaks (over 80 hours long). Participant P3 reflected on the issues with build time and coverage: “Yes, we have brought attention to these aspects in terms of build duration and code coverage. We are trying to understand this aspect in meetings, until this morning, the project managers asked me why the tests are taking so long”. This was a problem they were unable to resolve during the case study. This instability is also reflected in the high variation of “Build Health” and “Time to Fix a Broken Build” practices, as Participant P2 declares: “Last week we had some problems with infrastructure, we ended up having to do some configurations and such. Because of it, the build health metric started to have a lower value”. The “Commits per Day” tended to decrease in the last 4 weeks of the case study compared to the first 4 weeks. This trend can be attributed to the conclusion of the development cycle, during which the project shifted its focus toward documentation. As reinforced by participant P4: “… as everyone is focused on documentation, there shouldn’t be any more commits, except something from the README, something very specific like that.”
In Organization C, the project initially demonstrated good values for CI practices, which was surprising considering the self-declared BASIC level of CI in the initial survey. However, there was a significant degradation in all metrics during the break for the end-of-year holidays (the red dotted circle). Despite this, the project experienced a quick recovery after the holiday period for most of the practices, such as: “Build Health”, “Build Activity” and “Time to Fix a Broken Build”. The CI monitoring helped with this recovery after the holiday break as Participant 5 said: “There was also a period at the end of last year of the build being generated in a broken way, and as soon as we came back, one of the first things we did was fix it… We had already observed this in Gitlab, but the monitoring tool also contributed, in a way, complemented this need to fix it”
Organization A experimented with a notable enhancement in CI practices throughout the case study, particularly in coverage practices where reports indicate that monitoring played a crucial role in identifying and addressing areas that needed improvement. On the other hand, Organization B faced challenges due to the instability of its CI environment, impairing significant improvements in the project’s CI levels. Finally, in Organization C, the break for the end-of-year holiday has caused a temporary reduction in CI practices, which were quickly improved with the help of monitoring.
5. Discussion
In this section, we discuss the implications of our study.
During the conduction of the case study, it was observed that the automation of the build process and the execution of automated tests are the primary factors considered when defining the CI maturity level of projects. However, when CI monitoring was introduced, developers emphasized the monitoring process. Monitoring highlighted problems, served as a reminder, and motivated the team to enhance the CI process. They have a clear perception that this improvement could have an impact on project quality.
Combining CI practices into a single tool gives an overview of the CI process. It would be important for tools like GitLab to incorporate monitoring of CI practices. This integration would help in monitoring CI maturity. The call for tools like GitLab to incorporate monitoring of CI practices reflects a desire for more seamless and integrated solutions. Direct integration of CI monitoring into popular developer platforms, such as GitLab, streamlines the workflow for development teams. It supports efficient code collaboration and provides real-time insights into the health and effectiveness of CI processes. This integration aligns with the broader industry movement towards DevOps practices, where the collaboration between development and operations is tightly integrated.
Implications for researchers:
This case study showed that there are still gaps in monitoring development processes. Even though CI is already a widespread practice, few studies seek to evaluate its correct application. There is a field of study exploring new CI practices not evaluated in this study or even broader practices such as DevOps.
Implications for practitioners:
This study demonstrated that developers highly desire the monitoring of CI practices and derive significant benefits from its adoption. Therefore, commercial CI services like GitHub, GitLab, or Jenkins would greatly benefit from incorporating dashboards to monitor development practices such as CI practices. This addition would add substantial value to these tools and contribute to improving the quality of projects utilizing them.
6. Threats to Validity
In this section, we discuss the limitations of our study and how we attempted to mitigate them. Based on the work of (Elazhary et al., 2021), we utilized the framework outlined by (Roller and Lavrakas, 2015) to organize this section. This framework examines the sources of variability and possible bias in qualitative methods.
6.1. Data Gathering
This section focuses on discussing the accuracy and reliability of the collected data.
Scope: While our case study has provided valuable insights into Continuous Integration practices, the scope of our study was limited to three software development organizations, in our immediate geographic area, so we may have physical access and permission to install the monitoring suite and collect data. The projects were represented by 5 participants. As a result, the generalization of our findings may be constrained to the project’s context, and the participants can not express a whole project understanding of CI. We tried to select participants who had knowledge about the project and the project’s CI infrastructure, as a way to mitigate this problem.
Period of Case Study: The period in which the study was conducted - 8 weeks - may not be enough to have a complete effect on any action taken by the team after becoming aware of the CI project maturity. Furthermore, as we were immersed in the project’s environment, no developer was dedicated exclusively to the case study, continuing with their daily tasks, so they may not have had time to implement changes in the analyzed project, even though they were alerted by the monitoring of CI practices.
The case study was carried out from November 2023 until February 2024. This period coincides with the holiday season. Therefore, some projects were at the end of the development cycle, with lower demand, final tests or project documentation were carried out, which had a slight impact on some metrics during a certain project period. A break in interviews was held between 2023-12-22 and 2024-01-19 to mitigate this problem.
Changes in the Project’s Team: Organization A had changes in its development team during the case study, which may have influenced the evolution of CI practice values in RQ5. An intern left the Organization A at the beginning of the case study and three new interns entered at the end of the case study (these interns did not participate in the case study). In Organization B, one of the participants had a vacation during the last 2 weeks of the study, which may have influenced the data in RQ4 and RQ5.
6.2. Data Analysis
This section is dedicated to examining the accuracy of the analysis performed and the interpretations drawn from the study.
Thematic Analysis: In the Thematic Analysis, we acknowledge the potential for bias arising from the authors’ subjective interpretations of interview transcriptions. To mitigate researcher bias, the transcript coding was sent to each participant in the case study to validate the generated themes.
Transcription: To minimize errors during the transcription phase of the case study, we recorded all interviews. Subsequently, we transcribed them using a voice-to-text transcription tool [^2] https://www.veed.io/tools/audio-to-text. The transcripts were then verified by the first author, who listened to the interviews and compared them against the automatic tool transcription to ensure accuracy.
7. Conclusion
In this paper, we conducted a case study across three projects to investigate the benefits of monitoring CI practices in real-world scenarios. The monitoring of CI practices remains a research gap and is often overlooked by the industry. Our main objective was to address this gap in practical settings. To achieve this, we developed a monitoring suite tool deployed in the projects’ environments, providing study participants with the ability to monitor CI practices with minimal intrusion.
The findings from our study revealed that monitoring CI practices plays a crucial role in the software development lifecycle. It highlights potential issues within the CI process and provides an overview of the CI states, serving as a constant motivator and reminder for developers to enhance their CI processes. Our study indicates developers’ strong desire to integrate CI monitoring dashboards with popular CI services tools, such as GitLab, GitHub, and Jenkins, suggesting that seamless integration could further optimize and facilitate the development workflow.
In future work, it is highly recommended to expand the scope of CI monitoring practices to incorporate additional DevOps practices, such as metrics from the DORA (DevOps Research and Assessment)  [^3] https://dora.dev/. Extending the monitoring framework to include deployment and infrastructure metrics offers an extended perspective on the software delivery lifecycle. Including these broader metrics aligns with industry best practices and ensures a more robust evaluation of DevOps practices’ overall efficiency and effectiveness within a project.
Acknowledgements.
This work is partially supported by INES (www.ines.org.br), CNPq grant 465614/2014-0, CAPES grant 88887.136410/2017-00, and
FACEPE grants APQ-0399-1.03/17 and PRONEX APQ/0388-1.03/14.